Below is a purchasing schedule and a schedule showing the dollar amount of those purchases remaining...

  1. Below is a purchasing schedule and a schedule showing the dollar amount of those purchases remaining as an inventory balance for successive months.

    Purchases

    Feb

    Mar

    Apr

    May

    June

    Feb

    1650

    330

    174

    Mar

    2280

    456

    234

    Apr

    2880

    576

    302

    May

    3740

    748

    402

    June

    5190

    1038

    NA

    630

    810

    1050

    1440

    What is the balance fraction for March (the yellow highlighted cell)?

    a.

    10.26%

    b.

    10.55%

    c.

    24%

    d.

    20%

  2. Following Question 39, what is the June's balance fraction for purchases made in May (the green highlighted cell)?

    a.

    10.75%

    b.

    10.49%

    c.

    20%

    d.

    22%
Answer #1

Answer 1:

d. 20%

Explanation:

Balance fraction for March = 456 / 2280 = 20%

Hence option d is correct and other options a, b and c are incorrect.

Answer 2:

a. 10.75%

Explanation:

June's balance fraction for purchases made in May = 402 / 3740 = 10.75%

Hence option a is correct and other options b, c and d are incorrect.